# Bandit Sequential Posted Pricing via Half-Concavity

### v1

@@ 3.1 Proof Overview
Now we give the regret of Algorithm 3. 3.2 guarantees that for $i\geq 2$, every price vector we test in $\mathsf{Alg}_{i}$ has regret bounded by $\epsilon_{i-1}$. Therefore, the total regret of Algorithm 3 can be bounded by
@@
In this paper, we further assume that $n^{5}=o(\sqrt{T})$, otherwise an $O(\mathsf{poly}(n))$ regret algorithm is trivial. Then, $O(n^{5}\log^{2}T)$ can be upper bounded by $o(n^{2.5}\sqrt{T}\log T)$ and the regret bound of Algorithm 3 is $O(n^{2.5}\sqrt{T}\log T)$.
@@
3.3 Main Sub-Routine for Sequential Buyers: Proof of 3.2
@@
3.3.1 Notation
    The intuition of function $R_{i}(p)$ is: When finding $\hat{p}_{i}$, intuitively we hope to fix other prices. For a buyer $j$ before $i$, we set $p_{j}$ to be $r^{s}_{j}$ to maximize the probability of testing $i$. For buyer $k$ from suffix $S_{i+1}$, we set $p_{k}$ to be the settled $\hat{p}_{k}$ to minimize the error from the suffix. It should be noted that $R_{i}(p)$ can be also written as
    $R_{i}(p)~{}=~{}\mathsf{Pre}(S_{i})+P_{i}\big{(}p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p)\big{)}.$
    We will mainly use the second definition in our proofs.
    The intuition of $R^{*}_{i}(p)$ is : notice that $\arg\max_{p}r^{*}_{i}(p)=p^{*}_{i}$. A key idea of our algorithm is to approach approach $p^{*}_{i}$ via estimating $R^{*}_{i}(p)$. It should be noted that $R^{*}_{i}(p)$ can be also written as
    We will mainly use the second definition in our proofs.
3.3.2 Step 1: Finding $(\hat{p}_{1},\cdots,\hat{p}_{n})$ to Approximate $(p^{*}_{1},\cdots,p^{*}_{n})$
The first step of our subroutine is to find $(\hat{p}_{1},\cdots,\hat{p}_{n})$ that approximates $(p^{*}_{1},\cdots,p^{*}_{n})$. Specifically, we hope to achieve the following lemma:
@@ 3.3.2 Step 1: Finding $(\hat{p}_{1},\cdots,\hat{p}_{n})$ to Approximate $(p^{*}_{1},\cdots,p^{*}_{n})$
Given $\epsilon\geq\frac{1}{\sqrt{T}}$, intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ that satisfies $p^{*}_{i}\in[\ell_{i},r_{i}]$ for all $i\in[n]$, there exists an algorithm that tests $O(\frac{n^{5}\log^{2}T}{\epsilon^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in[\ell_{i},r^{s}_{i}]$ for all $i\in[n]$ and outputs $(\hat{p}_{1},\cdots,\hat{p}_{n})$. With probability $1-T^{-4}$, we have $P_{i}\cdot\mathsf{Loss}(S_{i})~{}\leq~{}5(n-i+1)\delta$ holds for all $i\in[n]$, where $\delta:=\frac{\epsilon}{100n^{2}}$ is the scaled error parameter.
3.3 is based on a key subroutine that finds $\hat{p}_{i}$ such that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})$ is small. The following lemma suggests that such a subroutine exists:
Lemma 3.4.
There exists a subroutine $\mathsf{SubR}$ that takes $R_{i}(p),[\ell_{i},r_{i}],\delta$ as input, tests $O(\delta^{-2}\log^{2}T)=O(\epsilon^{-2}n^{4}\log^{2}T)$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p\in[\ell_{i},r_{i}-{T^{-100}}],\hat{p}_{i+1},\cdots,\hat{p}_{n})$, and outputs $\hat{p}_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ that satisfies with probability $1-T^{-5}$ that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
The design of $\mathsf{SubR}$ requires the idea of generalized half-concavity. We first assume the existence of $\mathsf{SubR}$ and defer the detailed algorithm to Section 3.3.3.
With the help of 3.4, we are ready to give our algorithm and the proof of 3.3:
@@
3       Call $\mathsf{SubR}$ with input $R_{i}(p),[\ell_{i},r_{i}],\delta$ and get $\hat{p}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
@@
We show that Algorithm 4 is the algorithm we desire for 3.3. 3.4 guarantees that calling $\mathsf{SubR}$ one time runs $O(\frac{n^{4}\log^{2}T}{\epsilon^{2}})$ rounds. Therefore, Algorithm 4 tests $O(\frac{n^{5}\log^{2}T}{\epsilon^{2}})$ rounds in total. The constraint $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ is also guaranteed by 3.4.
It only remains to prove $P_{i}\cdot\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$. Notice that Algorithm 4 calls $\mathsf{SubR}$ $n$ times. Each call fails with probability $T^{-5}$. By the the union bound [^2] Recall that we assumed $n^{10}=o(T)$, and therefore $n<T$., with probability $1-T^{-4}$, $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds for all $i\in[n]$. Now we prove $P_{i}\cdot\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ assuming $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds.
@@
Case 2: $p^{*}_{i}\in[r^{s}_{i},r_{i}]$. For this case, we first prove the following claim:
Claim 3.5.
For all $i\in[n]$, we have $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})$.
Proof.
For $p\leq\mathsf{Opt}(S_{i+1})$, we have
$p\cdot(1-F_{i}(p))+\mathsf{Opt}(S_{i+1})\cdot F_{i}(p)~{}$ $\leq~{}\mathsf{Opt}(S_{i+1})\cdot(1-F_{i}(p))+\mathsf{Opt}(S_{i+1})\cdot F_{i}(p)$
$=~{}\mathsf{Opt}(S_{i+1})$
$=~{}\mathsf{Opt}(S_{i+1})\cdot(1-F_{i}(\mathsf{Opt}(S_{i+1})))+\mathsf{Opt}(S_{i+1})\cdot F_{i}(\mathsf{Opt}(S_{i+1})).$
Therefore, there must be
$p^{*}_{i}~{}:=~{}\arg\max_{p}\big{(}p\cdot(1-F_{i}(p))+\mathsf{Opt}(S_{i+1})\cdot F_{i}(p)\big{)}~{}\geq~{}\mathsf{Opt}(S_{i+1}).\qed$
3.5 helps prove the following inequality:
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}=p^{*}_{i},p^{*}_{i+1},\cdots,p^{*}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}=r^{s}_{i-1},p^{*}_{i+1},\cdots,p^{*}_{n})$
[…]
Here, the first inequality is from the fact that
$F_{i}(r^{s}_{i})\cdot(r^{s}_{i}-\mathsf{Opt}(S_{i+1}))~{}\leq~{}F_{i}(p^{*}_{i})\cdot(p^{*}_{i}-\mathsf{Opt}(S_{i+1})).$
This is true because we have $0\leq F_{i}(r^{s}_{i})\leq F_{i}(p^{*}_{i})$, $r^{s}_{i}-\mathsf{Opt}(S_{i+1})\leq p^{*}_{i}-\mathsf{Opt}(S_{i+1})$ and $p^{*}_{i}-\mathsf{Opt}(S_{i+1})\geq 0$. Then, for quantities $A,B,C,D$ satisfying $0\leq A\leq B,C\leq D,D\geq 0$, it is easy to verify that $A\cdot C\leq B\cdot D$. Now we are ready to give the induction step:
@@
3.3.3 Subroutine $\mathsf{SubR}$: Proof of 3.4
In this subsection, we introduce the sub-routine $\mathsf{SubR}$ that finds $\hat{p}_{i}$ satisfying $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. Recall that

### v2

@@ 3.2 Notation
Now we give the regret of Algorithm 4. 3.2 guarantees that for $i\geq 2$, every price vector we test in $\mathsf{Alg}_{i}$ has regret at most $\epsilon_{i-1}$. Hence, the total regret of Algorithm 4 can be bounded by
@@
In this paper, we consider the setting that $n$ is a fixed parameter while $T$ goes to infinity, and further assume that $n^{5}=o(\sqrt{T})$, otherwise an $O(\mathsf{poly}(n))$ regret algorithm is trivial. Then, $O(n^{5}\log^{2}T)$ can be upper bounded by $o(n^{2.5}\sqrt{T}\log T)$ and the regret bound of Algorithm 4 is $O(n^{2.5}\sqrt{T}\log T)$.
@@
3.4 Two-Steps Sub-Routine: Proof of 3.2
@@
As discussed in Section 3.1, the main sub-routine defined in 3.2 contains two steps: the first step is to find a group of approximately optimal prices $(\hat{p}_{1},\cdots,\hat{p}_{n})$, and the second step is to use $(\hat{p}_{1},\cdots,\hat{p}_{n})$ as a benchmark to update the upper and lower bounds via binary search. We first give corresponding key lemmas for these two steps.
Lemma 3.3 (Step 1).
Given $\delta:=\frac{\epsilon}{100n^{2}}\geq\frac{1}{\sqrt{T}}$, intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ that satisfies $p^{*}_{i}\in[\ell_{i},r_{i}]$ for all $i\in[n]$, there exists an algorithm that tests $O(\frac{n\log^{2}T}{\delta^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in[\ell_{i},r^{s}_{i}]$ for all $i\in[n]$ and outputs $(\hat{p}_{1},\cdots,\hat{p}_{n})$. With probability $1-T^{-4}$, we have the following for all $i\in[n]$:
  • $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
  • $P_{i}\cdot\mathsf{Loss}_{i}\leq 5(n-i+1)\delta$.
Lemma 3.4 (Step 2).
Given $\delta:=\frac{\epsilon}{100n^{2}}\geq\frac{1}{\sqrt{T}}$, intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ that satisfies $p^{*}_{i}\in[\ell_{i},r_{i}]$ for all $i\in[n]$, and near-optimal prices $\hat{p}_{1},\cdots,\hat{p}_{n}$ that satisfy the conditions in 3.3, there exists an algorithm that tests $O(\frac{n\log^{2}T}{\delta^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in[\ell_{i},r^{s}_{i}]$ for all $i\in[n]$ and outputs $[\ell^{\prime}_{1},r^{\prime}_{1}],\cdots,[\ell^{\prime}_{n},r^{\prime}_{n}]$. With probability $1-T^{-4}$, we have the following:
  • $[\ell^{\prime}_{i},r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$.
  • $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$.
    $R(p^{*}_{1},\cdots,p^{*}_{n})~{}-~{}R(p_{1},\cdots,p_{n})~{}\leq~{}\epsilon.$
Combining 3.3 and 3.4 immediately proves 3.2: we run the algorithms defined in 3.3 and 3.4 sequentially. Both algorithms use $O(\delta^{-2}\cdot n\log T)=O(\epsilon^{-2}\cdot n^{5}\log T)$ rounds, so the constraint on number of rounds in 3.2 is satisfied. The feasibility of each tested price vector is guaranteed by both 3.3 and 3.4. Finally, for the required conditions of the output, each algorithm succeeds with probability $1-T^{-4}$. By the union bound, with probability $1-2T^{-4}>1-T^{-3}$, both algorithms succeed simultaneously. Then, the statements in 3.2 is guaranteed by 3.4.
It only remains to prove 3.3 and 3.4. Next, we prove 3.3 in Section 3.5, and 3.4 in Section 3.6.
3.5 Step 1: Finding Near-Optimal $(\hat{p}_{1},\cdots,\hat{p}_{n})$ to Prove 3.3
In this section, we give the algorithm defined in 3.3. Our algorithm is based on a key sub-algorithm that finds $\hat{p}_{i}$ such that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. The following lemma states the existence of such a sub-algorithm:
@@
There exists a sub-algorithm $\mathsf{SubAlg}$ that takes hidden function $R_{i}(p)$, interval $[\ell_{i},r_{i}]$, and parameter $\delta$ as input, tests $O(\delta^{-2}\log^{2}T)$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p\in[\ell_{i},r^{s}_{i}],$ $\hat{p}_{i+1},\cdots,\hat{p}_{n})$, and outputs $\hat{p}_{i}\in[\ell_{i},r^{s}_{i}]$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ with probability $1-T^{-5}$.
The design of $\mathsf{SubAlg}$ requires the idea of generalized half-concavity discussed in Section 3.1. Since the proof requires some involved technical details, we defer it separately to Section 3.7. With the help of 3.5, we give our algorithm and the proof of 3.3.
@@
3       Call $\mathsf{SubAlg}$ with input $R_{i}(p),[\ell_{i},r_{i}],\delta$ and get $\hat{p}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
@@
We show that Algorithm 5 is the desired algorithm for 3.3. 3.5 guarantees that calling $\mathsf{SubAlg}$ one time runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds. Therefore, Algorithm 5 tests $O(\frac{n\log^{2}T}{\delta^{2}})$ rounds in total. For the succeed probability, Algorithm 5 calls $\mathsf{SubAlg}$ $n$ times. Each call fails with probability $T^{-5}$. By the the union bound [^2] Recall that we assumed $n=o(T)$., with probability $1-T^{-4}$, all calls of $\mathsf{SubAlg}$ in Algorithm 5 succeed. Then, the constraints $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ and $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ are directly guaranteed by 3.5.
It only remains to prove $P_{i}\cdot\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$. Now we prove this statement assuming $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds.
@@
Case 2: $p^{*}_{i}\in[r^{s}_{i},r_{i}]$. For this case, we have
@@
where the first inequality is from the following fact: let $C:=R(1,\cdots,1,p_{i}=1,p^{*}_{i+1},\cdots,p^{*}_{n})$ be the optimal revenue from suffix $S_{i}$, we have
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},r^{s}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})$
[…]
where the second inequality is from the fact that $p^{*}_{i}\geq C$, because $p^{*}_{i}$ is the maximizer of function $R^{*}_{i}(p)$, and $R^{*}_{i}(p^{*}_{i})-R^{*}_{i}(C)=P_{i}\cdot(1-F_{i}(p^{*}_{i}))\cdot(p^{*}_{i}-C)\geq 0$ implies $p^{*}_{i}\geq C$.
@@
3.6 Step 2: Determining New Intervals via Binary Search to Prove 3.4
Claim 3.6.
Proof.
With 3.6, the idea of our binary search algorithm is clear: We first set $R_{i}(\hat{p}_{i})$ be the benchmark. For a price $p$, we can estimate $R_{i}(p)$ and compare $R_{i}(p)+5(n-i)\delta$ with $R_{i}(\hat{p}_{i})$. If $R_{i}(p)+5(n-i)\delta<R_{i}(\hat{p}_{i})$, it is evident to say $R^{*}_{i}(p)<R^{*}_{i}(\hat{p}_{i})$, and therefore $p$ can’t be $p^{*}_{i}$. Otherwise, we keep $p$ as a candidate of $p^{*}_{i}$. Since function $R_{i}(p)$ is single-peaked, the experiment above can be done via two binary search algorithms, and extra steps are needed to take care of the errors from the uncertainty - we only have estimates $\hat{R}_{i}(p)$ and $\hat{R}_{i}(\hat{p}_{i})$, instead of $R_{i}(p)$ and $R_{i}(\hat{p}_{i})$. Now, we state the following lemma that achieves our idea:
Lemma 3.7.
Assume 3.3 holds. Given hidden function $R_{i}(p)$, interval $[\ell_{i},r_{i}]\ni p^{*}_{i}$, scaled $\delta:=\frac{\epsilon}{100n^{2}}>1/\sqrt{T}$, price $\hat{p}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$, there exists a binary search algorithm that runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}\in[\ell_{i},r^{s}_{i}],\hat{p}_{i+1},\cdots,\hat{p}_{n})$ and outputs with probability $1-T^{-5}$ $[\ell^{\prime}_{i},r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$ that satisfies:
  • $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$.
To prove 3.7, we first give the following Algorithm 6, and then show Algorithm 6 is the desired algorithm for 3.7.
2 Let $\ell_{b1}\leftarrow\ell_{i},r_{b1}\leftarrow\hat{p}_{i}$.
3 while $r_{b1}-\ell_{b1}\geq{T^{-100}}$ do
4       Let $m=\frac{\ell_{b1}+r_{b1}}{2}$.
7
  // Second binary search to determine $r^{\prime}_{i}$
11 if $\hat{R}_{i}(r-{T^{-100}})\geq\hat{R}_{i}(\hat{p}_{i})-2\delta$ then
       // Special Case: original upper bound is good
12       Let $r^{\prime}_{i}\leftarrow r_{i}$
13else
14       Let $\ell_{b2}\leftarrow\hat{p}_{i},r_{b2}\leftarrow r_{i}-{T^{-100}}$.
15       while $r_{b2}-\ell_{b2}\geq{T^{-100}}$ do
16             Let $m=\frac{\ell_{b2}+r_{b2}}{2}$.
19
20      Let $r^{\prime}_{i}\leftarrow r_{b2}$.
Output: $[\ell^{\prime}_{i},r^{\prime}_{i}]$.
ALGORITHM 6 Getting $[\ell^{\prime}_{i},r^{\prime}_{i}]$ via Binary Search
Proof of 3.7.
We show Algorithm 6 satisfies the statements in 3.7. Our proof starts from verifying Algorithm 6 runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds with $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$. For the number of tested rounds, Algorithm 6 runs two standard binary searches. In one binary search, the subroutine tests one price with $C\cdot\frac{\log T}{\delta^{2}}$ rounds, and the binary search stops when the length of the interval is less than ${T^{-100}}$. Therefore, The total number of rounds is $C\cdot\frac{\log T}{\delta^{2}}\cdot O(\log T^{100})=O(\frac{\log^{2}T}{\epsilon^{2}})$. For the range constraint of all tested prices, Algorithm 6 directly guarantees that every tested price falls in $[\ell_{i},r_{i}-{T^{-100}}]$.
Claim 3.8.
Proof.
Now we show $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$. The key idea is to use 3.6. For clarity we first restate 3.6: See 3.6
Now we prove $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$. We first show that $p^{*}_{i}\geq\ell_{b1}$. Here, $\ell_{b1}$ refers to the value after the while loop is finished. $p^{*}_{i}\geq\ell_{b1}$ holds because $\ell_{b1}$ is updated to be at least $m$ when we observe $\hat{R}_{i}(m)<\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$. Combining this with 3.8 implies $R_{i}(m)<R_{i}(\hat{p}_{i})-5(n-i)\delta$, i.e., $m\notin I^{\prime}$. Since $R_{i}(p)$ is single-peaked, it’s sufficient to update $\ell_{b1}$ to at least $m$ and $I^{\prime}\subseteq[\ell_{b1},1]$ holds. Symmetrically, a similar argument gives $p^{*}_{i}\leq r_{b2}=r^{\prime}_{i}$.
It remains to show $p^{*}_{i}\geq\ell^{\prime}_{i}$. $\ell^{\prime}_{i}$ and $\ell_{b1}$ are different when we have $\hat{R}(\ell_{b1})<\hat{R}(r_{b1})-3\delta$. Combining this inequality with 3.8 gives $R(r_{b1})-R(\ell_{b1})>\delta>r_{b1}-\ell_{b1}$. This is in contrast to the constrained Lipschitzness from 3.9. Then, there must be $r_{b1}<\mathsf{Rev}(S_{i+1})<\mathsf{Opt}(S_{i+1})\leq p^{*}_{i}$, where the last inequality $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})$ is implied by the fact that $R^{*}_{i}(p^{*}_{i})-R^{*}_{i}(\mathsf{Opt}(S_{i+1}))=P_{i}\cdot(1-F_{i}(p^{*}_{i}))\cdot(p^{*}_{i}-\mathsf{Opt}(S_{i+1}))\geq 0$. Therefore, it’s feasible to update $\ell^{\prime}_{i}$ to $r_{b1}$ in this case.
We first prove that $R_{i}(\hat{p}_{i})-R(\ell^{\prime}_{i})\leq 5(n-i)\delta+11\delta$. The binary search subroutine guarantees that $\hat{R}_{i}(r_{b1})\geq\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$ after the binary search loop is finished. Then, 3.8 ensures that $R_{i}(\hat{p}_{i})-R_{i}(r_{b1})\leq 5(n-i)\delta+4\delta$. Then, if $\hat{R}_{i}(\ell_{b1})<\hat{R}_{i}(r_{b1})-3\delta$, we set $\ell^{\prime}_{i}$ to be $r_{b1}$. Otherwise, we have $\hat{R}_{i}(\ell_{b1})\geq\hat{R}_{i}(r_{b1})-3\delta$ and $\ell^{\prime}_{i}$ is set to be $\ell_{b1}$. Then,
[…]
Otherwise, the constrained Lipschitzness of $R_{i}(p)$ gives
Finally, we end this section proving 3.4 via 3.7.
Proof of 3.4.
The algorithm for 3.4 is to use 3.7 $n$ times and get $n$ new confidence intervals. 3.7 succeeds with probability $1-T^{-5}$. By the union bound, with probability $1-n\cdot T^{-5}>1-T^{-4}$, our algorithm for 3.4 succeeds, and the conditions $[\ell^{\prime}_{i}r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$ and $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$ are guaranteed by 3.7. It only remains to show that each price vector in the new intervals has regret no more than $\epsilon$.
We instead prove the following statement via induction, assuming the statements in 3.3 and 3.7 hold: For $p_{i},\cdots,p_{n}$ that satisfies $p_{j}\in[\ell^{\prime}_{j},r^{\prime}_{j}-{T^{-100}}]$ for all $j\in\{i,\cdots,n\}$, we have
The base case is $i=n$. 3.7 gives
For the induction step, we hope to show
under the induction hypothesis
Now we finish the induction step:
[…]
$+10(n-i+1)\delta+14\delta+50(n-i)^{2}\delta$
[…]
Here, the first inequality is from the inequality $P_{i}\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ in 3.3. The second inequality is from 3.3 that guarantees $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. The third inequality is from 3.7 that guarantees $R_{i}(\hat{p}^{*}_{i})-R_{i}(p_{i})\leq 5(n-i)\delta+15\delta$. The fourth inequality is from the fact that prices $p^{*}_{i+1},\cdots,p^{*}_{n}$ are optimal for suffix $S_{i+1}$. The fifth inequality is from the corollary of the induction hypothesis. Comparing the start and the end of the above inequality finishes the induction step.
Finally, taking $i=1$ with the fact that $\delta=\frac{\epsilon}{100n^{2}}$ finishes the proof of 3.4. ∎
3.7 Sub-Algorithm $\mathsf{SubAlg}$: Proof of Lemma 3.5
We now introduce the sub-algorithm $\mathsf{SubAlg}$ that finds $\hat{p}_{i}$ satisfying $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$, and prove the missing 3.5 that will complete the proof of Step 1. We restate the lemma for convenience.
See 3.5
We start from introducing the following new notations that we need in this subsection:
With the help of these new notations, we can rewrite $R_{i}(p)$ as